OpenStudy (anonymous):

2. How did velocity (v) and velocity squared (v2) relate to the height from which the ball was released? Use the relationships from your graphs in your explanation. Which one, if any, was linearly related to the height? This problem is from a lab about conservation of energy. I need some help. Also, this is my first time using Open Study, so I am not sure how this work.

OpenStudy (irishboy123):

energy conservation is often the easiest way to navigate this \(\frac{1}{2} m v^2 = mgh\) if something has fallen a distance h in the earth's gravitational field, it has gained kinetic energy \(\frac{1}{2} m v^2\) and surrendred potenial energy \( mgh\) that is just an opening observation. any qu's, just ask 😋

OpenStudy (anonymous):

Thank you for the response! So, the ball with the velocity squared has a greater amount on kinetic energy? I guess that would mean the ball that is higher has more kinetic energy and less potential energy, while it is in motion? @IrishBoy123 ... Also, how can you determine which is linear to the height?

OpenStudy (irishboy123):

\(v^2\) should be linearly realted to the height

OpenStudy (anonymous):

Thank you for the help @IrishBoy123 .. I really appreciate it.

OpenStudy (irishboy123):

mon plaisir 💥

OpenStudy (anonymous):

Merci, tu sont tres sympa! (I think that is right, I am a little rusty on French) @IrishBoy123 .. Anyway thanks again!

OpenStudy (irishboy123):

moi aussi !! 🤔

OpenStudy (anonymous):

k